OpenStudy (anonymous):

I don't get how to do this? Can someone help please! In the following example of "Simplifying an Expression", what reason would be used for step C? Simplify the expression 3k + 7 + 2k and justify each step 3k + 7 + 2k = 3k + 2k + 7 A = (3k + 2k) + 7 B = (3 + 2)k + 7 C = 5k + 7 D A. Addition B. Commutative Property C. Associative Property D. Distributive Property

OpenStudy (anonymous):

You are using the Distributive Property "backwards". Usually, the Distributive Property is stated as: A(B + C) = AB + AC, but in this case you are going the other direction. AB + AC = A(B + C), or in this particular case 3k + 2k = (3 + 2)k.